# CharCom: Composable Identity Control for Multi-Character Story Illustration

### v1

@@ 1. Introduction
Text-to-image (T2I) diffusion models have shown remarkable progress in generating photorealistic and semantically rich images from textual prompts. Mainstream diffusion models (Rombach et al., 2022; Gu et al., 2023b) are now widely used in creative domains such as visual storytelling, educational illustration, and children’s books. However, despite their ability to generate high-resolution and visually vivid images, these models still struggle to maintain consistent character identity across multiple generated scenes (Avrahami et al., 2023; Zhou et al., 2024).
@@
This consistency is particularly critical for storybook illustration, where readers expect not only high visual quality but also temporal and semantic continuity of characters throughout a story. Current pipelines often suffer from identity drift, where the same character appears with different facial features, hairstyles, clothing, or poses across scenes. Such inconsistencies disrupt narrative coherence and undermine the practical utility of diffusion models in real-world storytelling applications (Maharana et al., 2022; Pan et al., 2022). To mitigate character inconsistency, early approaches (Ruiz et al., 2023) perform full model fine-tuning for each subject to encode identity features, achieving high fidelity but incurring heavy computational overhead and poor scalability. More recently, IP-Adapter (Ye et al., 2023) introduces a plug-and-play visual adapter that projects identity features into the model’s latent space. While efficient and training-free, it remains limited in complex narrative scenarios, particularly those involving multiple interacting characters or dynamic scene transitions (Wang et al., 2024; Wu et al., 2024). These limitations highlight the central challenge of our work: Ensuring consistent visual identity for each character across a sequence of story scenes.
@@
In this work, we propose CharCom, a lightweight, scalable, and composable framework for character-consistent story illustration using diffusion models. Inspired by Low-Rank Adaptation (LoRA) (Hu et al., 2021), our method represents each character with an independently trained rank-4 identity adapter. These adapters can be composed at inference time without modifying or retraining the backbone diffusion model, enabling multi-character scene generation with minimal training data and no fine-tuning overhead. To further enhance narrative coherence, we design a structured prompt template that explicitly controls key elements such as character roles, emotions, actions, and background context (Cheng et al., 2024; Tao et al., 2024). As shown in Figure 1, CharCom achieves more consistent character rendering and scene coherence than several popular methods. Visually, facial features remain highly stable, and characters retain their distinctive appearance across multiple scenes, offering a more faithful and engaging visual narrative.
@@
We further introduce a character-aware evaluation framework based on GPT-4o (OpenAI, 2024), which performs multimodal visual question answering (VQA) to assess identity fidelity and temporal coherence. Compared to conventional CLIP-based metrics (Radford et al., 2021a), this approach offers more semantically grounded and human-aligned assessments (Huang et al., 2023). All experiments are conducted on Shakoo Maku [^1] https://www.shakoomaku.com/, an Arabic-language animated IP for children aged 3–6 featuring recurring characters across scenes. To support evaluation, we construct a small-scale synthetic benchmark comprising 20 storylines, each with 5 prompts containing overlapping characters, generated with GPT-4o (Hurst et al., 2024) to ensure narrative diversity. CharCom demonstrates clear advantages over strong existing methods, achieving consistent generation with minimal computational overhead. Overall, our main contributions can be summarized as follows:
@@ 2. Related Work
Text-to-image (T2I) diffusion models have advanced rapidly following the introduction of Denoising Diffusion Probabilistic Models (DDPMs) (Ho et al., 2020), which generate images by iteratively denoising Gaussian noise. DDIM (Song et al., 2021) further improves inference speed via a non-Markovian sampling process.
Latent diffusion models (Rombach et al., 2022; Gu et al., 2023a) further improve efficiency by leveraging variational autoencoders (VAEs) to operate in latent space, significantly reducing memory and compute cost while enabling high-resolution generation. These models commonly adopt CLIP (Radford et al., 2021a) embeddings and classifier-free guidance to enhance prompt fidelity. FLUX (Kang et al., 2025) introduces dual encoders and flow-matching loss for better prompt alignment under constrained compute budgets. In this work, we adopt FLUX as our base model and focus on enabling character consistency on top of this efficient backbone.
@@ 2.1. Character Identity Preservation
Maintaining consistent character identity across diverse prompts is a core challenge for diffusion-based models. Earlier works like DreamBooth (Ruiz et al., 2023) achieve high fidelity by fine-tuning the entire diffusion backbone with subject-specific prompts, but suffer from high computational cost and poor scalability. Textual Inversion (Wei et al., 2025) offers a more efficient alternative by optimizing subject-specific token embeddings while freezing the model backbone, though it often compromises visual fidelity and controllability.
To reduce fine-tuning overhead, Low-Rank Adaptation (LoRA) (Hu et al., 2021; Kumari et al., 2022) introduces a parameter-efficient strategy by decomposing weight updates into low-rank matrices. This method has been extended to support multi-concept blending (He and Yao, 2025; Wu et al., 2025), though these approaches typically require offline merging or scene-level planning to avoid interference among identities. Parallel to these, adapter-based inference methods such as IP-Adapter (Ye et al., 2023) and InstantID (Wang et al., 2024) offer training-free alternatives by injecting visual features through attention or facial embeddings. However, their conditioning mechanisms are often underconstrained in multi-character scenarios, leading to identity entanglement and reduced consistency.
@@ 2.2. Multi-Character and Story-Level Modeling
Recent efforts have extended T2I generation toward story-level consistency and multi-character coherence. OneActor (Xie et al., 2023) uses pseudo-labels and compositional training to improve generalization, but is limited to single-subject scenarios. ORACLE (Wu et al., 2024) enhances identity fidelity within LoRA-based generation through mutual information constraints, though this increases training complexity and limits scalability. More comprehensive systems such as StoryMaker (Zhou et al., 2024) and CharacterFactory (Wang et al., 2025) aim to model entire narratives by integrating modules like scene decomposition, character disentanglement, or GAN-based identity embeddings.
@@ 4. Methodology
We propose CharCom, a modular, composable framework for identity-consistent multi-character, multi-scene story illustration, built on top of FLUX (Kang et al., 2025), a recent diffusion-based text-to-image model. Unlike static conditioning methods such as IP-Adapter (Ye et al., 2023) or mutual-information-based LoRA systems like ORACLE (Wu et al., 2024), CharCom requires neither joint optimization nor backbone modifications. It trains lightweight, per-character adapters independently and composes them dynamically at inference, enabling flexible multi-character control without retraining. Formally, the model learns $\mathcal{G}_{\Theta}(p,S)$ to generate images that are temporally and semantically consistent given scene prompts and character identities. As illustrated in Figure 2, CharCom runs atop the FLUX diffusion backbone, with per-character identity adapters selectively injected into attention and multi-layer perceptron layers at inference to enable controllable composition.
Figure 2. Overview of CharCom architecture built on FLUX (Kang et al., 2025). CLIP-L and T5-XXL encode prompts via dual encoders, and a frozen VAE handles image latents. Character-specific LoRA adapters are trained independently and composed at inference.
@@ 4.1. System Pipeline
  • Prompt Encoding: We employ a dual-encoder setup with CLIP-L (Radford et al., 2021b) and T5-XXL (Chung, 2022) to jointly encode prompts $\mathcal{P}$ into complementary visual and semantic representations, enhancing grounding and alignment.
  • Latent Denoising: A DiT-style U-Net denoiser is guided by FlowMatch (Lipman et al., 2022) and a temporal consistency loss, enforcing coherent transitions across scenes.
@@ 4.2. Low-Rank Adapter Training and Fusion
We adopt low-rank adaptation (Hu et al., 2021) to efficiently encode character-specific identity features into the diffusion model parameters. For any weight matrix $W\in\mathbb{R}^{d\times d}$ in FLUX backbone, we insert a learnable low-rank residual:
@@ 4.3. Consistency-Oriented Prompt Design
where each component encodes distinct semantic factors to minimize cross-character interference. Here, $\tau_{c}$ denotes a unique character-specific token (e.g., Shakoo Maku Lulu) used to explicitly identify each individual character. Each character is associated with an independently trained adapter, which is linked to its corresponding trigger token. Embedding these tokens directly into the prompt provides explicit identity grounding and enables the model to selectively activate the corresponding identity modules during inference. $\Phi_{c}$ encodes fixed, character-specific attributes such as age, hairstyle, and clothing. We utilize Qwen2.5-Max (Team, 2024) to generate high-quality textual descriptions for each character that appears in Shakoo Maku in advance. These descriptions are compressed into 15–25 token sequences via $\Gamma(\cdot)$, balancing semantic richness with token efficiency. This compact representation ensures efficient processing within the attention mechanism. $\mathcal{A}$ encodes task-specific actions (e.g., “sits beside”, “reads a book”) extracted from the input sentence, guiding the character’s pose and interactions. In contrast, $\mathcal{S}$ captures the global rendering style, including aesthetic cues such as the color palette and illustration type.
@@ 5.1. Experiment Setup
We compare CharCom against three baselines: (i) Vanilla FLUX, the original diffusion model without character conditioning, (ii) FLUX + IP-Adapter, which incorporates image-based personalization, and (iii) Static-All, a prompt-unaware LoRA baseline inspired by ORACLE (Wu et al., 2024) that loads all character adapters with equal weights ($w_{c}=1$) at inference. CharCom is evaluated in compositional inference mode without additional fine-tuning. This setting reveals potential interference that occurs when identity modules are injected without semantic disentanglement.

### v2

@@ 1. Introduction
Text-to-image (T2I) diffusion models have shown remarkable progress in generating photorealistic and semantically rich images from textual prompts. Mainstream diffusion models (rombach2022highresolution; gu2023assessing) are now widely used in creative domains such as visual storytelling, educational illustration, and children’s books. However, despite their ability to generate high-resolution and visually vivid images, these models still struggle to maintain consistent character identity across multiple generated scenes (avrahami2023roleconsistency; zhou2024storymaker).
@@
This consistency is particularly critical for storybook illustration, where readers expect not only high visual quality but also temporal and semantic continuity of characters throughout a story. Current pipelines often suffer from identity drift, where the same character appears with different facial features, hairstyles, clothing, or poses across scenes. Such inconsistencies disrupt narrative coherence and undermine the practical utility of diffusion models in real-world storytelling applications (maharana2022storydalleap; pan2022flintstonessv). To mitigate character inconsistency, early approaches (ruiz2023dreambooth) perform full model fine-tuning for each subject to encode identity features, achieving high fidelity but incurring heavy computational overhead and poor scalability. More recently, IP-Adapter (ye2023ipadapter) introduces a plug-and-play visual adapter that projects identity features into the model’s latent space. While efficient and training-free, it remains limited in complex narrative scenarios, particularly those involving multiple interacting characters or dynamic scene transitions (wang2024instantid; oracle2024oracle). These limitations highlight the central challenge of our work: Ensuring consistent visual identity for each character across a sequence of story scenes.
@@
In this work, we propose CharCom, a lightweight, scalable, and composable framework for character-consistent story illustration using diffusion models. Inspired by Low-Rank Adaptation (LoRA) (hu2021lora), our method represents each character with an independently trained rank-4 identity adapter. These adapters can be composed at inference time without modifying or retraining the backbone diffusion model, enabling multi-character scene generation with minimal training data and no fine-tuning overhead. To further enhance narrative coherence, we design a structured prompt template that explicitly controls key elements such as character roles, emotions, actions, and background context (cheng2024autostudiocc; tao2024coinal). As shown in Figure 1, CharCom achieves more consistent character rendering and scene coherence than several popular methods. Visually, facial features remain highly stable, and characters retain their distinctive appearance across multiple scenes, offering a more faithful and engaging visual narrative.
@@
We further introduce a character-aware evaluation framework based on GPT-4o (openai2024gpt4o), which performs multimodal visual question answering (VQA) to assess identity fidelity and temporal coherence. Compared to conventional CLIP-based metrics (radford2021clip), this approach offers more semantically grounded and human-aligned assessments (huang2023vqascore). All experiments are conducted on Shakoo Maku [^1] https://www.shakoomaku.com/, an Arabic-language animated IP for children aged 3–6 featuring recurring characters across scenes. To support evaluation, we construct a small-scale synthetic benchmark comprising 20 storylines, each with 5 prompts containing overlapping characters, generated with GPT-4o (hurst2024gpt) to ensure narrative diversity. CharCom demonstrates clear advantages over strong existing methods, achieving consistent generation with minimal computational overhead. Overall, our main contributions can be summarized as follows:
@@ 2. Related Work
Text-to-image (T2I) diffusion models have advanced rapidly following the introduction of Denoising Diffusion Probabilistic Models (DDPMs) (ho2020ddpm), which generate images by iteratively denoising Gaussian noise. DDIM (song2020ddim) further improves inference speed via a non-Markovian sampling process.
Latent diffusion models (rombach2022highresolution; gu2023ass) further improve efficiency by leveraging variational autoencoders (VAEs) to operate in latent space, significantly reducing memory and compute cost while enabling high-resolution generation. These models commonly adopt CLIP (radford2021clip) embeddings and classifier-free guidance to enhance prompt fidelity. FLUX (kang2025fluxak) introduces dual encoders and flow-matching loss for better prompt alignment under constrained compute budgets. In this work, we adopt FLUX as our base model and focus on enabling character consistency on top of this efficient backbone.
@@ 2.1. Character Identity Preservation
Maintaining consistent character identity across diverse prompts is a core challenge for diffusion-based models. Earlier works like DreamBooth (ruiz2023dreambooth) achieve high fidelity by fine-tuning the entire diffusion backbone with subject-specific prompts, but suffer from high computational cost and poor scalability. Textual Inversion (wei2025personalizedig) offers a more efficient alternative by optimizing subject-specific token embeddings while freezing the model backbone, though it often compromises visual fidelity and controllability.
To reduce fine-tuning overhead, Low-Rank Adaptation (LoRA) (hu2021lora; kumari2022multiconceptco) introduces a parameter-efficient strategy by decomposing weight updates into low-rank matrices. This method has been extended to support multi-concept blending (he2025conceptrolcc; wu2025lesstomoregu), though these approaches typically require offline merging or scene-level planning to avoid interference among identities. Parallel to these, adapter-based inference methods such as IP-Adapter (ye2023ipadapter) and InstantID (wang2024instantid) offer training-free alternatives by injecting visual features through attention or facial embeddings. However, their conditioning mechanisms are often underconstrained in multi-character scenarios, leading to identity entanglement and reduced consistency.
@@ 2.2. Multi-Character and Story-Level Modeling
Recent efforts have extended T2I generation toward story-level consistency and multi-character coherence. OneActor (oneactor2023oneactor) uses pseudo-labels and compositional training to improve generalization, but is limited to single-subject scenarios. ORACLE (oracle2024oracle) enhances identity fidelity within LoRA-based generation through mutual information constraints, though this increases training complexity and limits scalability. More comprehensive systems such as StoryMaker (zhou2024storymaker) and CharacterFactory (wang2025characterfactory) aim to model entire narratives by integrating modules like scene decomposition, character disentanglement, or GAN-based identity embeddings.
@@ 4. Methodology
We propose CharCom, a modular, composable framework for identity-consistent multi-character, multi-scene story illustration, built on top of FLUX (kang2025fluxak), a recent diffusion-based text-to-image model. Unlike static conditioning methods such as IP-Adapter (ye2023ipadapter) or mutual-information-based LoRA systems like ORACLE (oracle2024oracle), CharCom requires neither joint optimization nor backbone modifications. It trains lightweight, per-character adapters independently and composes them dynamically at inference, enabling flexible multi-character control without retraining. Formally, the model learns $\mathcal{G}_{\Theta}(p,S)$ to generate images that are temporally and semantically consistent given scene prompts and character identities. As illustrated in Figure 2, CharCom runs atop the FLUX diffusion backbone, with per-character identity adapters selectively injected into attention and multi-layer perceptron layers at inference to enable controllable composition.
Figure 2. Overview of CharCom architecture built on FLUX (kang2025fluxak). CLIP-L and T5-XXL encode prompts via dual encoders, and a frozen VAE handles image latents. Character-specific LoRA adapters are trained independently and composed at inference.
@@ 4.1. System Pipeline
  • Prompt Encoding: We employ a dual-encoder setup with CLIP-L (radford2021learning) and T5-XXL (https://doi.org/10.48550/arxiv.2210.11416) to jointly encode prompts $\mathcal{P}$ into complementary visual and semantic representations, enhancing grounding and alignment.
  • Latent Denoising: A DiT-style U-Net denoiser is guided by FlowMatch (lipman2022flow) and a temporal consistency loss, enforcing coherent transitions across scenes.
@@ 4.2. Low-Rank Adapter Training and Fusion
We adopt low-rank adaptation (hu2021lora) to efficiently encode character-specific identity features into the diffusion model parameters. For any weight matrix $W\in\mathbb{R}^{d\times d}$ in FLUX backbone, we insert a learnable low-rank residual:
@@ 4.3. Consistency-Oriented Prompt Design
where each component encodes distinct semantic factors to minimize cross-character interference. Here, $\tau_{c}$ denotes a unique character-specific token (e.g., Shakoo Maku Lulu) used to explicitly identify each individual character. Each character is associated with an independently trained adapter, which is linked to its corresponding trigger token. Embedding these tokens directly into the prompt provides explicit identity grounding and enables the model to selectively activate the corresponding identity modules during inference. $\Phi_{c}$ encodes fixed, character-specific attributes such as age, hairstyle, and clothing. We utilize Qwen2.5-Max (qwen25) to generate high-quality textual descriptions for each character that appears in Shakoo Maku in advance. These descriptions are compressed into 15–25 token sequences via $\Gamma(\cdot)$, balancing semantic richness with token efficiency. This compact representation ensures efficient processing within the attention mechanism. $\mathcal{A}$ encodes task-specific actions (e.g., “sits beside”, “reads a book”) extracted from the input sentence, guiding the character’s pose and interactions. In contrast, $\mathcal{S}$ captures the global rendering style, including aesthetic cues such as the color palette and illustration type.
@@ 5.1. Experiment Setup
We compare CharCom against three baselines: (i) Vanilla FLUX, the original diffusion model without character conditioning, (ii) FLUX + IP-Adapter, which incorporates image-based personalization, and (iii) Static-All, a prompt-unaware LoRA baseline inspired by ORACLE (oracle2024oracle) that loads all character adapters with equal weights ($w_{c}=1$) at inference. CharCom is evaluated in compositional inference mode without additional fine-tuning. This setting reveals potential interference that occurs when identity modules are injected without semantic disentanglement.